OpenStudy (anonymous):

Line perpendicular to y= -4x+6

ganeshie8 (ganeshie8):

in xy plane, there will be many lines perpendicular to a given line

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

Do you mind explaining a little bit more? I am still lost. @ganeshie8

ganeshie8 (ganeshie8):

the question looks incomplete.. are you really sure you're not hiding anything ?

ganeshie8 (ganeshie8):

usually they give a point through which the perpendicular line goes so that you pin down the exact line

OpenStudy (anonymous):

No there is nothing more to the question, but there are options

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

what are the options ?

OpenStudy (anonymous):

y=16x-24 -4y+x=6 -4y-x=6 -4x=-4x+2

OpenStudy (anonymous):

@ganeshie8